I need help answering t test stastic equation .

2. After receiving the state tax assessment of $87,500 for his house, John decided to appeal this assessment. He was told that his assessment was made as the average of all houses in his neighborhood. John hired a consulting company to randomly sample and assess eight houses in the neighborhood. The results are: $85,200; $86,400; $84,250; $85,300; $87,600; $84,750; $90,100 and $85,800. The consultant verified that house values in John's neighborhood are normally distributed. Use a 0.05 level of significance and test the hypothesis that the average assessed value of all houses in John's neighborhood is less than the $87,500 he was originally assessed. Did John get the state to lower its tax assessment of his house?
Answer #1

From the given data, the calculation are as follows:

# X Mean (x - mean)2
1 85200 86175 950625
2 86400 86175 50625
3 84250 86175 3705625
4 85300 86175 765625
5 87600 86175 2030625
6 84750 86175 2030625
7 90100 86175 15405625
8 85800 86175 140625
n 8
Sum 689400
Average 86175.00
SS 25080000
Variance = SS/n-1 3582857.143
Std Dev=Sqrt(Variance) 1892.84367

\mu = $87,500, \bar{x} = $86,175, s = $1892.843, n = 8

The Hypothesis:

H0: \mu = $87,500 : The average assessed value in Johns neighbourhood is equal to $87,500.

Ha: \mu < $87500 : The average assessed value in Johns neighbourhood is less than $87,500.

This is a left tailed test

The Test Statistic: Since the population standard deviation is unknown, we use the students t test.

The test statistic is given by the equation:

t = \frac{\bar{x}-\mu }{\frac{s}{\sqrt{n}}} = \frac{86175-87500}{\frac{1892.843}{\sqrt{8}}} = -1.98

t observed = -1.98

The p Value:    The p value (Left Tail) for t = -1.98, for degrees of freedom (df) = n-1 = 7, is; p value = 0.0440

The Critical Value:   The critical value (Left Tail) at \alpha = 0.05, for df = 7, tcritical = -1.8946

The Decision Rule:   If tobserved is < -1.8946, Then Reject H0.

Also if P value is < \alpha (0.05), Then Reject H0.

The Decision: Since tobserved (-1.98) is < -1.89 , We Reject H0.

Also since P value (0.0440) is < 0.05, We Reject H0.

The Conclusion: Reject H0. There is sufficient evidence at the 95% significance level to conclude that the average assessed value in Johns neighbourhood is less than $87,500.

Therefore Yes, John got the state to lower its tax assessment of his house.